OpenStudy (anonymous):

Use graphical approximation techniques or an equation solver to approximate the desired interest rate. A person makes annual payments of $1000 into an ordinary annuity. At the end of 5 years, the amount in the annuity is 5767.98. What annual nominal compounding rate has this annuity earned? Rate needs to be rounded 2 decimal places.

OpenStudy (anonymous):

\[x^5 = 767.98\]

OpenStudy (amistre64):

the present value of an annuity should be 0; also refered to as an internal rate of return i think

OpenStudy (amistre64):

1000/(1+r)^1 1000/(1+r)^2 1000/(1+r)^3 1000/(1+r)^4 1000/(1+r)^5 ------------- sum = 5767.98

OpenStudy (amistre64):

lol, almost had that right ....

OpenStudy (amistre64):

sum would have to be zero for net present value :) best thing to do according to my accounting classes was to guess at it

OpenStudy (amistre64):

\[FV=PV\frac{1-(1+r)^n}{r}\] \[5767.98=1000\frac{1-(1+r)^5}{r}\] \[5.76798=\frac{1-(1+r)^5}{r}\] \[5.76798r=1-(1+r)^5\] see where the 2 graphs intersect